given that \( x \) is a random variable from a binomial distribution with parameter \( (n, p) \), Using the m.g.f technique show that the expected value and variance of \( x \) is \( n p \) and \( n p q \) respectively

1
Expert's answer
2022-01-24T02:46:32-0500

For Binomial distribution

"f(x)=\binom{n}{x}p^xq^{n-x}\\
M(t)=E(e^{tx})=e^{tx}\sum_{x=0}^n\binom{n}{x}p^xq^{n-x}\\
=\sum_{x=0}^n\binom{n}{x}(pe^t)^xq^{n-x}\\
=(pe^t+q)^n\\
M'(t)=npe^t(pe^t+q)^{n-1}\\
M'(0)=E(x)=np(p+q)^{n-1}=np\\
M''(t)=npe^t(pe^t+q)^{n-2}(n-1)pe^t+npe^t(pe^t+q)^{n-1}
\\
M''(0)=np(p+q)^{n-2}(n-1)p+np(p+q)^{n-1}\\
M''(0)=np(n-1)p+np=n(n-1)p^2+np\\
var(x)=M''(0)-(M''(0))^2\\
=n(n-1)p^2+np-(np)^2\\
=np-np^2=np(n-p)=npq"